OpenStudy (anonymous):

a right triangle has one leg of length 8 and another leg of length 9. What is the hypotenuse of the triangle?

OpenStudy (anonymous):

\[\sqrt{145}\]

OpenStudy (angela210793):

Hypotenuse=sqr(8^2+9^2)=sqr(64+81)=sqr145 Pythagorean theorem